# Hairy black holes in extended Einstein-Maxwell-scalar theories with magnetic charge and kinetic couplings

### v1

@@ IV Concrete models and numerical analyses
Figs. 3-4 show the behavior of $|h-f|$ for the several values of $c_{1}$, $c_{2}$, or $c_{3}$. The numerical results support the analytic solutions given in Eqs. (4.3)-(4.5). From Eq. (4.3), there is a threshold at which the dominant contribution switches from the linear term to the quadratic term. For increasing $c_{1}$, our numerical result in Fig. 3 shows that the metric difference grows first, but it starts to be suppressed around $c_{1}\gtrsim{\cal O}(100)$. In terms of large $c_{2}$, $|h-f|$ increases monotonically in Fig. 3. This result shows good agreement with Eq. (4.4). However, $c_{2}$ cannot be large without the limit since the upper bound (4.6) exists. For a too large $c_{2}$ to violate the condition (4.6) with fixed $P$, numerical integration tends to be unstable. Indeed, we numerically confirmed that the singular behavior starts to appear around $c_{2}=1870$, for the model parameters $\mu=0.1$, $P=0.001M_{\rm pl}r_{h}$, $\phi_{0}=0.1M_{\rm pl}$. We note that the large $c_{2}$ limit has been discussed in Ref. Lee and Weinberg (1991) under the assumption of the weak gravitational field. It has been analytically shown that the solutions around the horizon ($r\gtrsim r_{h}$) reduce to the RN metric possessing only a magnetic charge in such a case. In contrast to the previous study, we do not assume the weak field approximation and find that the solutions can deviate from the RN metric in large $c_{2}$ but singular behavior occurs at some point. When taking large $c_{3}$, Eq. (4.5) implies that the metric difference converges to finite values. We numerically confirmed that this saturation realizes around $c_{3}\gtrsim{\cal O}(100)$ as shown in Fig. 4. Hence, we can distinguish the three couplings by $|h-f|$ in the large coupling limit.
@@
$\frac{|h-f|}{(r-r_{h})^{2}}$ $\simeq\left|-\frac{\left[\mu M_{\rm pl}^{3}r_{h}^{2}-P^{2}(c_{1}\phi_{0}+M_{\rm pl})\right]c_{1}M_{\rm pl}}{(1-\mu)(c_{1}\phi_{0}+M_{\rm pl})\,d_{1}\phi_{0}r_{h}^{2}}+\frac{M_{\rm pl}^{6}r_{h}^{2}}{(1-\mu)\,d_{1}^{2}\phi_{0}^{2}}\right|+\mathcal{O}\left(\frac{1}{d_{1}}\right)$ $\left(|d_{1}|\gg\{|d_{2}|,\,|d_{3}|\}\right)\,,$ (4.11)
$\frac{|h-f|}{(r-r_{h})^{2}}$ $\simeq\left|-\frac{Pc_{2}M_{\rm pl}\sqrt{2\mu M_{\rm pl}^{2}r_{h}^{2}-P^{2}}}{(1-\mu)\,d_{2}\phi_{0}r_{h}^{2}}+\frac{M_{\rm pl}^{6}r_{h}^{2}}{(1-\mu)\,d_{2}^{2}\phi_{0}^{2}}\right|+\mathcal{O}\left(\frac{1}{d_{2}}\right)$ $\left(|d_{2}|\gg\{|d_{1}|,\,|d_{3}|\}\right)\,,$ (4.12)
$\frac{|h-f|}{(r-r_{h})^{2}}$ $\simeq\left|-\frac{(2\mu M_{\rm pl}^{2}r_{h}^{2})c_{3}M_{\rm pl}^{2}}{2(1-\mu)(c_{3}\phi_{0}+M_{\rm pl})\,d_{3}\phi_{0}r_{h}^{2}}+\frac{M_{\rm pl}^{6}r_{h}^{2}}{(1-\mu)\,d_{3}^{2}\phi_{0}^{2}}\right|+\mathcal{O}\left(\frac{1}{d_{3}}\right)$ $\left(|d_{3}|\gg\{|d_{1}|,\,|d_{2}|\}\right)\,,$ (4.13)
@@ V Conclusions
In this paper, we focused on the hairy BH solutions electrically and magnetically charged on the static and spherically symmetric background based on the quadratic interaction $f_{2}(\phi,X,F,\tilde{F},Y)$ in the $U(1)$ gauge-invariant SVT theories. The hairy BH solutions originated from the cubic and the quartic interactions, $\mathcal{L}_{3}$ and $\mathcal{L}_{4}$, in the $U(1)$ gauge-invariant SVT theories in Refs. Heisenberg et al. (2018) in the absence of magnetic charge. It would be of interest to investigate whether the presence of magnetic charge gives rise to a new type of hairy BH solution in the full $U(1)$ gauge-invariant SVT theories. It is also of interest to step forward to BH perturbation analyses to examine the stability of our solutions. The BH perturbations within the theory including scalar-vector coupling were studied in Refs. Heisenberg et al. (2018); Gannouji and Baez (2022); Kase and Tsujikawa (2023) in the absence of magnetic charge. It is necessary to clarify how the existence of magnetic charge affects the dynamics of perturbations as well as stability conditions.
@@ Acknowledgements
RK is supported by the Grant-in-Aid for Early-Career Scientists of the JSPS No. 20K14471 and the Grant-in-Aid for Scientific Research (C) of the JSPS No. 23K034210.

### v2

@@ IV.2 Large coupling limit of each interaction and possible parameter space
Figs. 3-4 show the behavior of $|h-f|$ for the several values of $c_{1}$, $c_{2}$, or $c_{3}$. The numerical results support the analytic solutions given in Eqs. (4.3)-(4.5). From Eq. (4.3), there is a threshold at which the dominant contribution switches from the linear term to the quadratic term. For increasing $c_{1}$, our numerical result in Fig. 3 shows that the metric difference grows first, but it starts to be suppressed around $c_{1}\gtrsim{\cal O}(100)$. In terms of large $c_{2}$, $|h-f|$ increases monotonically in Fig. 3. This result shows good agreement with Eq. (4.4). However, $c_{2}$ cannot be large without the limit since the upper bound (4.6) exists. For a too large $c_{2}$ to violate the condition (4.6) with fixed $P$, numerical integration tends to be unstable. Indeed, we numerically confirmed that the singular behavior starts to appear around $c_{2}=1870$, for the model parameters $\mu=0.1$, $P=0.001M_{\rm pl}r_{h}$, $\phi_{0}=0.1M_{\rm pl}$. The violation of the condition (4.6) implies that the resultant spacetime significantly deviates from RN spacetime. Since the asymptotic solutions (3.3)-(3.8) around the horizon are obtained under the assumption that the couplings behave as corrections to the RN solution, they cannot approximate the solutions violating the condition (4.6). We note that the large $c_{2}$ limit has been discussed in Ref. Lee and Weinberg (1991) under the assumption of the weak gravitational field. It has been analytically shown that the solutions around the horizon ($r\gtrsim r_{h}$) reduce to the RN metric possessing only a magnetic charge in such a case. In contrast to the previous study, we do not assume the weak field approximation and find that the solutions can deviate from the RN metric in large $c_{2}$ but singular behavior occurs at some point. When taking large $c_{3}$, Eq. (4.5) implies that the metric difference converges to finite values. We numerically confirmed that this saturation realizes around $c_{3}\gtrsim{\cal O}(100)$ as shown in Fig. 4. Hence, we can distinguish the three couplings by $|h-f|$ in the large coupling limit.
@@ IV.3 Effects of $X$-dependence in $g_{i}(\phi,X)$ on scalar hair
$\frac{|h-f|}{(r-r_{h})^{2}}$ $\simeq\left|-\frac{\left[\mu M_{\rm pl}^{3}r_{h}^{2}-P^{2}(c_{1}\phi_{0}+M_{\rm pl})\right]c_{1}M_{\rm pl}}{(1-\mu)(c_{1}\phi_{0}+M_{\rm pl})\,d_{1}\phi_{0}r_{h}^{2}}+\frac{M_{\rm pl}^{6}r_{h}^{2}}{(1-\mu)\,d_{1}^{2}\phi_{0}^{2}}\right|+\mathcal{O}\left(\frac{1}{d_{1}^{3}}\right)$ $\left(|d_{1}|\gg\{|d_{2}|,\,|d_{3}|\}\right)\,,$ (4.11)
$\frac{|h-f|}{(r-r_{h})^{2}}$ $\simeq\left|-\frac{Pc_{2}M_{\rm pl}\sqrt{2\mu M_{\rm pl}^{2}r_{h}^{2}-P^{2}}}{(1-\mu)\,d_{2}\phi_{0}r_{h}^{2}}+\frac{M_{\rm pl}^{6}r_{h}^{2}}{(1-\mu)\,d_{2}^{2}\phi_{0}^{2}}\right|+\mathcal{O}\left(\frac{1}{d_{2}^{3}}\right)$ $\left(|d_{2}|\gg\{|d_{1}|,\,|d_{3}|\}\right)\,,$ (4.12)
$\frac{|h-f|}{(r-r_{h})^{2}}$ $\simeq\left|-\frac{(2\mu M_{\rm pl}^{2}r_{h}^{2})c_{3}M_{\rm pl}^{2}}{2(1-\mu)(c_{3}\phi_{0}+M_{\rm pl})\,d_{3}\phi_{0}r_{h}^{2}}+\frac{M_{\rm pl}^{6}r_{h}^{2}}{(1-\mu)\,d_{3}^{2}\phi_{0}^{2}}\right|+\mathcal{O}\left(\frac{1}{d_{3}^{3}}\right)$ $\left(|d_{3}|\gg\{|d_{1}|,\,|d_{2}|\}\right)\,,$ (4.13)
@@ V Conclusions
In this paper, we focused on the hairy BH solutions electrically and magnetically charged on the static and spherically symmetric background based on the quadratic interaction $f_{2}(\phi,X,F,\tilde{F},Y)$ in the $U(1)$ gauge-invariant SVT theories. The hairy BH solutions originated from the cubic and the quartic interactions, $\mathcal{L}_{3}$ and $\mathcal{L}_{4}$, in the $U(1)$ gauge-invariant SVT theories were studied in Refs. Heisenberg and Tsujikawa (2018) in the absence of magnetic charge. It would be of interest to investigate whether the presence of magnetic charge gives rise to a new type of hairy BH solution in the full $U(1)$ gauge-invariant SVT theories. It is also of interest to step forward to BH perturbation analyses to examine the stability of our solutions. The BH perturbations within the theory including scalar-vector coupling were studied in Refs. Heisenberg et al. (2018); Gannouji and Baez (2022); Kase and Tsujikawa (2023); Zhang and Kase (2024) in the absence of magnetic charge. It is necessary to clarify how the existence of magnetic charge affects the dynamics of perturbations Nomura et al. (2020); Chen et al. (2024) as well as stability conditions in the $U(1)$ gauge-invariant SVT theories..
@@ Acknowledgements
R. K. is supported by the Grant-in-Aid for Early-Career Scientists of the JSPS No. 20K14471 and the Grant-in-Aid for Scientific Research (C) of the JSPS No. 23K03421.